OpenStudy (mckenzieandjesus):

Find the missing coefficient. (4d-4)(3d-2)= 12d^2 + blank d+8

OpenStudy (mckenzieandjesus):

4 –20 –4 20

OpenStudy (arivillafana):

LHS = (4d-4)(3d-2) = (4d)(3d) + (4d)(-2) + (-4)(3d) + (-4)(-2) = 12d^2 – 8d – 12d +8 = 12d^2 – 20d + 8 ------ (1) But RHS = 12d^2 + _d + 8 ---------(2) Therefore from (1) and (2), Missing Coefficient is -20.

OpenStudy (jdoe0001):

hint: expand and simplify the left-hand-side

OpenStudy (calculusxy):

Use the FOIL method.

OpenStudy (mckenzieandjesus):

\[\left( 4d-4 \right)\left( 3d-2 \right)=12d^2+[blank]d + 8\]

OpenStudy (arivillafana):

I just gave her the answer and everyone is ignoring it lol

OpenStudy (jdoe0001):

yes, so expand and simplify the left-hand-side it'd end up in a quadratic, and you'd see the middle term :)

OpenStudy (calculusxy):

\[(4d-4)(3d-2) = 4d(3d) \times 4d(-2) \times -4(3d) \times -4(-2)\]

OpenStudy (calculusxy):

I am pretty sure that this is the foil method right?

OpenStudy (calculusxy):



OpenStudy (mckenzieandjesus):

what do u mean by expand?

OpenStudy (mckenzieandjesus):

oh okay

OpenStudy (mckenzieandjesus):

Thank you guys!

OpenStudy (calculusxy):

Did you get your answer?

OpenStudy (mckenzieandjesus):

Yes well @arivillafana gave it to me

OpenStudy (mckenzieandjesus):

Could u help me with this question? Match the expression with its name. 3x3 – 3x + 3 quadratic binomial cubic trinomial fourth-degree monomial not a polynomial

OpenStudy (arivillafana):

cubic

OpenStudy (mckenzieandjesus):

Are u sure?

OpenStudy (mckenzieandjesus):

I have a great grade right now & just dont wanna loose it haha

OpenStudy (arivillafana):

that's what i think the expression has 3 terms a leading term with a power of 3 is a cubic

OpenStudy (mckenzieandjesus):

Oh okay ur right

OpenStudy (mckenzieandjesus):

@calculusxy when i calculated that i came up with 132?